# Professor Leonardo P. Chamorro Professor Cristian Ricardo Constante Amores Professor Jie Feng

### v1

@@ 4.1 Introduction
Figure 4.1: (a) Vector plot of the field ($\dot{r}_{c},\dot{\alpha}$) for a Rankine vortex with $l/R=1.5$, $R$ is the size of the vortex core. The red point represents a neutrally stable fixed point (one end of the dumbbell is fixed at the vortex center). The orange and green lines denote the unstable and stable boundaries of an ’attracting set’. (b) Trajectory of the center of mass of the dumbbell in the Lamb-Oseen vortex for $r_{c}(0)=0.3$, $\phi_{c}(0)=\pi/4$, $\alpha(0)=0$. (c) The same as in (b) for $r_{c}(0)=1$, $\alpha(0)=-\pi/4$. Modified from (yerasi2022spirographic, ).
@@ B.2 Particle motion in symmetric Moffatt eddy flow with wall interactions
Figure B.3: The eddy streamline pattern in the latter case, from the symmetric stream function Eq. (2.1). Particles ($a_{p}=0.1$) follow closed trajectories (colored) for different particle initial positions.
@@ Appendix C Derivation of wall-parallel shear drag coefficient
To get the shear coefficient of the particle parallel velocity $f(\Delta)$, we start with the form from (pasol2011motion, ):
@@
According to (goldman1967slow2, ) as $\Delta\to\infty$, $f(\Delta)\simeq 1-\frac{5}{16}\Delta^{-3}$, which means that the order $\Delta^{-1}$ and $\Delta^{-2}$ must vanish, resulting in
@@
According to (williams1994particle, ) as $\Delta\to 0$, $f(\Delta)\simeq 1-\frac{1}{0.66-0.269\log(\frac{\Delta}{1+\Delta})}$. By matching all the parameters, we obtain $a=0.269$ and
@@ Appendix D Hydrodynamic resistance coefficients for wall interaction for wall-normal motion
The scalar quantities $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$ and $\mathcal{D}$ are dimensionless hydrodynamic resistances depending on $\Delta$. The analytical expressions for $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$ and $\mathcal{D}$ are given (as infinite sums) in (rallabandi2017hydrodynamic, ). For large separations($\Delta\gg 1$), one obtains four hydrodynamic resistances at leading order:

### v2

@@ 4.1 Introduction
Figure 4.1: (a) Vector plot of the field ($\dot{r}_{c},\dot{\alpha}$) for a Rankine vortex with $l/R=1.5$, $R$ is the size of the vortex core. The red point represents a neutrally stable fixed point (one end of the dumbbell is fixed at the vortex center). The orange and green lines denote the unstable and stable boundaries of an ’attracting set’. (b) Trajectory of the center of mass of the dumbbell in the Lamb-Oseen vortex for $r_{c}(0)=0.3$, $\phi_{c}(0)=\pi/4$, $\alpha(0)=0$. (c) The same as in (b) for $r_{c}(0)=1$, $\alpha(0)=-\pi/4$. Modified from (yerasi2022spirographic).
@@ B.2 Particle motion in symmetric Moffatt eddy flow with wall interactions
Figure B.3: The eddy streamline pattern in the latter case, from the symmetric stream function Eq. (2.1). Particles ($a_{p}=0.1$) follow closed trajectories (colored) for different particle initial positions. Adapted from Liu2024permanent.
@@ Appendix C Derivation of wall-parallel shear drag coefficient
To get the shear coefficient of the particle parallel velocity $f(\Delta)$, we start with the form from (pasol2011motion):
@@
According to (goldman1967slow2) as $\Delta\to\infty$, $f(\Delta)\simeq 1-\frac{5}{16}\Delta^{-3}$, which means that the order $\Delta^{-1}$ and $\Delta^{-2}$ must vanish, resulting in
@@
According to (williams1994particle) as $\Delta\to 0$, $f(\Delta)\simeq 1-\frac{1}{0.66-0.269\log(\frac{\Delta}{1+\Delta})}$. By matching all the parameters, we obtain $a=0.269$ and
@@ Appendix D Hydrodynamic resistance coefficients for wall interaction for wall-normal motion
The scalar quantities $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$ and $\mathcal{D}$ are dimensionless hydrodynamic resistances depending on $\Delta$. The analytical expressions for $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$ and $\mathcal{D}$ are given (as infinite sums) in (rallabandi2017hydrodynamic). For large separations($\Delta\gg 1$), one obtains four hydrodynamic resistances at leading order: